• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 19:08:49 on September 20, 2011

    Entry number 359169

    keyword=HRS DAQ Configurations and Status

    Here is a summary of the left and right HRS DAQ status.
    The script "startcoda" has been modified to correctly start the left or 
    right DAQ depending on wherther the script is run from adaql2 or adaql1,
    Left arm:
    1) Runs on adaql2 as user a-onl.  
    2) Currently consists of ROC3, ROC4 and TS11.
    3) Triggers T3 and T8 work, but need to check T4.
    3) Retiming module has been removed for repair, since the retimed L1A has
    a double pulse.
    4) Issue with the EB2 seg faulting may still need investigation.
    5) Currently taking cosmic data for checkout of the detectors.
    Right arm:
    1) Runs on adaql1 as user a-onl.
    2) Currently consists of ROC1, ROC2 and TS0.
    3) Trigger T1 and T8 exist, but need to check T2.
    4) Tomorrow Ryan, Melissa and Toby will check the timing for the ADC
    gates, then cosmic checkout will begin.

    A copy of this log entry has been emailed to: camsonne,rom,rbziel,kalyan,melissac,tbadman